# Ziyang Luo1 Nian Liu2 Fahad Shahbaz Khan2 Junwei Han1

### v1

@@ 1 Introduction
Humans perceive and interact with a complex, dynamic world by seamlessly integrating information from multiple sensory modalities. While vision plays a dominant role, auditory and textual cues are often crucial for accurately locating and comprehending objects of interest. This inherent multimodal nature of our perception highlights the necessity of developing systems that can achieve similar integrated understanding. Recent advancements in multimodal large language models, such as Qwen-Omni (Xu et al. 2025), and VideoLLaMA2 (Cheng et al. 2024), have demonstrated significant progress in audio-visual comprehension. However, tasks like Reference Audio-Visual Segmentation (Ref-AVS) (Wang et al. 2024b) still remain challenging, requiring models to precisely segment specific sounding objects by integrating textual references, audio cues, and visual information. This demands sophisticated multimodal reasoning alongside accurate spatial localization capabilities.
@@
While previous works have explored fusion methods (Wang et al. 2024b, 2025b; Radman and Laaksonen 2025) to align modalities, often employing an auxiliary text encoder with self-attention fusion, it remains unclear whether these models truly grasp the meaning of textual references or merely exhibit language bias. The “black box” nature of these approaches hinders genuine semantic understanding. To address these interpretability concerns, recent approaches have begun incorporating Large Language Models (LLMs) to enhance reasoning capabilities in Ref-AVS tasks, as demonstrated by Crab (Du et al. 2025). However, their existing implementations often rely on supervised fine-tuning with pre-defined simple reasoning templates. This approach tends to lead the LLM towards “memorizing” these fixed patterns, which could result in outputs that reflect post-hoc rationalization rather than genuine and systematic reasoning. Moreover, the design of jointly training complex language reasoning with precise visual segmentation introduces new challenges, which risks compromising the pixel-level precision of the original segmentation models as pointed out by (Liu et al. 2025).
@@
In this paper, we propose AURORA, a novel method that enhances both language comprehension and reasoning capabilities for audio-visual segmentation. Specifically, we leverage VideoLLaMA2 (Cheng et al. 2024) as our MLLM component to work in tandem with SAM (Kirillov et al. 2023) for segmentation, enabling reasoning-guided segmentation. To prevent models from merely rationalizing the final output, we introduce a structured Chain-of-Thought (CoT) (Wei et al. 2022) prompting mechanism via open-source Qwen-Omni (Xu et al. 2025) to cost-effectively generate diverse reasoning paths for training. It involves distinct analytical steps focusing sequentially on visual, audio, and textual reference cues, followed by a final integration of these analyses to derive the answer. To mitigate the potential conflict between reasoning and segmentation during joint training, we introduce a segmentation feature distillation loss. This loss distills knowledge from a segmentation-only model into the joint model, enabling it to acquire CoT reasoning capabilities without compromising segmentation performance. To endow the model with genuine reasoning ability, we further adopt a two-stage training strategy after supervised fine-tuning (SFT) with CoT, combining reflective learning and reinforcement learning via Group Reward Policy Optimization (GRPO) (Guo et al. 2025). In the corrective reflective-style stage, we transform low-quality reasoning paths from SFT-trained model into “reflective reasoning paths” by juxtaposing the initial flawed reasoning with a self-correction prompt followed by a Gemini-assisted high-quality revision. In the GRPO stage, we design a hybrid reward function incorporating format reward, IoU reward, and class reward to guide the model toward more robust and multimodally grounded reasoning.
@@ 2.1 Audio-Visual Segmentation
Audio-Visual Segmentation (AVS) aims to generate pixel-level masks for sounding objects and has attracted increasing research attention (Zhou et al. 2023). While some prior works have incorporated text modality into AVS, they primarily focus on class alignment without fully exploiting the semantic richness of textual information (Bhosale et al. 2023, 2024; Luo et al. 2025). Furthermore, traditional AVS methods (Zhou et al. 2023; Li et al. 2023a; Gao et al. 2024; Li et al. 2024; Yang et al. 2024; Ling et al. 2024) lack explicit guidance mechanisms, making it challenging to identify specific objects of interest within complex audio-visual scenes. To address this limitation, Ref-AVS (Wang et al. 2024b) introduces a reference-based framework that provides continuous segmentation guidance through audio-visual-text fusion via a multi-modal cue aggregation module. Building upon this approach, SAM2-LOVE (Wang et al. 2025b) further enhances the framework by integrating SAM2 with multimodal fusion, token propagation, and accumulation strategies. TSAM (Radman and Laaksonen 2025) enhances SAM with temporal modeling capabilities for spatio-temporal learning across video frames and replaces interactive prompting with data-driven prompts, thereby extending SAM’s functionality to dynamic video content. However, these methods treat text as an auxiliary modality and may fail to fully understand reference semantics.
Our work distinguishes itself by proposing a novel segmentation-centric framework integrating SAM (Kirillov et al. 2023) with VideoLLaMA2 (Cheng et al. 2024) for reference-guided segmentation, employing CoT (Wei et al. 2022) reasoning to incrementally decompose references and a segmentation feature distillation loss to preserve segmentation accuracy in SFT training.
@@ 2.2 Multimodal Language Models in Audio-Visual Scenes
In audio-visual understanding, predominant Multimodal Language Models (MLLMs) such as Qwen-Omni (Xu et al. 2025), and VideoLLaMA2 (Cheng et al. 2024) integrate audio, visual, and textual modalities to achieve comprehensive scene understanding. MEERKAT (Chowdhury et al. 2024) advances this field by constructing a fine-grained, large-scale audio-visual instruction-tuning dataset that endows models with temporal and spatial grounding capabilities. More specifically for reference AVS, CRAB (Du et al. 2025) attempts to unify multimodal understanding and segmentation within an LLM framework. However, their reliance on supervised fine-tuning with pre-defined reasoning templates may lead to superficial pattern memorization rather than fostering genuine systematic reasoning or deep semantic comprehension, limiting their robustness in complex scenarios. To address these limitations, following SFT, we introduce an additional two-stage training process. The first stage, corrective reflective-style training, encourages the model to refine its prior knowledge and calibrate its foundational perceptual abilities. The second stage, GRPO, facilitates robust self-improvement by optimizing the model’s reasoning pathways based on targeted reward functions.
@@ 3 Methodology
In this work, we introduce AURORA, a novel reasoning-enhanced framework for reference audio-visual segmentation that addresses the fundamental challenge of achieving genuine multimodal reasoning and maintaining precise segmentation performance. We implement this framework by integrating SAM (Kirillov et al. 2023) with VideoLLaMA2 (Cheng et al. 2024). Our approach employs a three-stage training procedure. In the first stage, we incorporate CoT reasoning using Qwen-Omni (Xu et al. 2025) for SFT with segmentation feature distillation loss. In the second stage, we enhance the model’s foundational accuracy through a novel error-correction stage using structured “reflective examples” that address common perceptual and knowledge errors. The third stage employs GRPO (Guo et al. 2025) to jointly enhance both reasoning capabilities and segmentation performance. Figure 1 illustrates the overall architecture of our proposed framework.
@@ 3.1 Stage1: SFT Training with CoT
Specifically, we first generate CoT answers using the open-source Qwen-Omni (Xu et al. 2025) model to conserve computational costs compared to proprietary models, and establish the initial format through carefully designed prompts. Through prompt engineering, we decompose the reference reasoning into four distinct steps as shown below.
@@
where $D$ denotes the fine-tuning dataset, and $\pi_{\theta}$ is the MLLM parameterized by $\theta$. Since Ref-AVS is a binary segmentation task, we follow the LISA framework (Lai et al. 2024) and append the token “It is [SEG]” after the CoT reasoning. The hidden state feature of the [SEG] token is then used as a visual prompt for the SAM decoder.
@@
While CoT reasoning has demonstrated effectiveness in language tasks, it presents a challenge for multi-modal segmentation tasks. We hypothesize that fine-tuning with a joint objective may compromise the pixel precision of the segmentation models (Liu et al. 2025), particularly when the language-based reasoning loss dominates the optimization dynamics. This, in turn, can degrade the representation of the [SEG] token, which is critical for triggering high-quality mask generation, thereby impairing segmentation performance. To address this issue and effectively decouple the optimization of reasoning and segmentation capabilities, we introduce a segmentation feature distillation loss as shown in the gray block in Figure 2.
@@ 3.2 Stage2: Corrective Reflective-Style Training
To endow our model with a more robust reasoning capability, we require high-quality examples of ideal thinking processes. We leverage a more powerful MLLM, Gemini, to serve as an expert annotator, whose role is to generate gold-standard reasoning paths by correcting flawed CoT samples. Specifically, we first prompt Gemini to critique the reasoning generated by our Stage 1 model. This process is focused on challenging samples from the training set, which we identify as those with both an IoU score below 0.6 and incorrect reasoning. This strict criterion ensures that corrective training targets only clear-cut failures where both the reasoning process and segmentation outcome are flawed. Adhering to the minimal modification principle (Yu et al. 2024), Gemini is instructed to rectify any identified flaws—such as incorrect conclusions or superficial analysis of audio information—by making the edits (e.g., modifying, adding, or deleting words). We denote the initial, potentially flawed CoT from the Stage 1 model as $y_{wrong}$ and the corrected version from Gemini as $y_{correct}$. We then construct a “reflective path” as follows:
@@
[…]
@@
AVSBench (Zhou et al. 2022) 23.2 51.1 37.2 32.4 54.7 43.5 27.8 52.9 40.3
AVSegFormer (Gao et al. 2024) 33.5 47.0 40.2 36.1 50.1 43.1 34.8 48.6 41.7
GAVS (Wang et al. 2024a) 28.9 49.8 39.4 29.8 49.7 39.8 29.4 49.8 39.6
@@
EEMC (Wang et al. 2024b) 34.2 51.3 42.8 49.5 64.8 57.2 41.9 58.1 50.0
SAM-LAVS (Wang et al. 2025b) 43.5 51.9 47.7 66.5 72.3 69.4 55.0 62.1 58.5
TSAM (Radman and Laaksonen 2025) 43.4 56.8 50.1 54.6 66.4 60.5 49.0 61.6 55.3
@@ 3.3 Stage3: Reinforcement Learning Training
Following the SFT stage, we incorporate reinforcement learning (RL) to further enhance the reasoning capability of the MLLM. Specifically, we adopt the GRPO (Guo et al. 2025) framework to enable self-refinement based on the relative quality of generated outputs within a group. However, GRPO was originally designed for text generation and does not directly align with the segmentation nature of Ref-AVS. To address this, we redesign the reward function tailored for segmentation as shown in the green block of Figure 2.
@@
where ${r}_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\text{old}}(o_{i,t}|q,o_{i,<t})}$. Following (Yu et al. 2025), we set the KL divergence coefficient $\beta$ to zero, effectively removing this term for simplicity.
@@ 4.1 Experimental Settings
We evaluated our method on the Ref-AVS benchmark dataset (Wang et al. 2024b), which contains 4,000 videos with manual pixel-level annotations and expressions. The dataset is divided into a training set (2,908 videos), a validation set (276 videos), and a test set (818 videos). The test set is further split into three subsets: (1) seen split: containing the same 39 categories as the training set; (2) unseen split: with 13 additional categories not present during training; (3) null split: testing cases where expressions refer to non-existent or invisible objects. As in (Wang et al. 2025b), we omit the null subset from evaluation.
@@
Our training unfolds in three stages. First, we perform an initial SFT for 720 steps using CoT reasoning generated by Qwen-Omni (Xu et al. 2025). In the second stage, we conduct a 300-step corrective fine-tuning phase. For this, we curate 1,505 reflective-style examples by using Gemini 2.5-Pro to refine incorrect outputs from the first stage. These are then mixed with 3,500 standard CoT samples to prevent catastrophic forgetting. The final stage consists of 500 steps of GRPO (Guo et al. 2025) training, where we generate 3 candidate responses per input for preference alignment. We use a batch size of 2 with 1 sample per device and gradient accumulation steps of 4. All experiments are conducted on two A40 GPUs.
@@
In line with the evaluation protocol of Ref-AVS (Wang et al. 2024b), we employ the Jaccard index ($\mathcal{J}$), F-score ($\mathcal{F}$), and their average ($&\mathcal{J}\&\mathcal{F}$) as the primary evaluation metrics.
@@ 4.2 Main Results
We conduct a comprehensive performance comparison of our proposed AURORA against a suite of top-performing SOTA methods on the Ref-AVS benchmark. These include three audio-based segmentation methods (Zhou et al. 2022; Gao et al. 2024; Wang et al. 2024a), two visual-based segmentation methods (Wu et al. 2022; Li et al. 2023b), three multi-modal methods (Wang et al. 2024b, 2025b; Radman and Laaksonen 2025), and one foundation-based method (Du et al. 2025). The detailed results, presented in Table 1, demonstrate that AURORA achieves new state-of-the-art performance across multiple key metrics.
On the seen test split, our model surpasses the second-best performing method by substantial margins. We attribute this significant leap in performance to our framework’s core design: unlike prior methods that rely on simple modality fusion (Wang et al. 2024b, 2025b; Radman and Laaksonen 2025) or train segmentation models from scratch (Du et al. 2025), AURORA leverages the power of a large language model for deep semantic reasoning and integrates it with a pre-trained foundation model for segmentation. More importantly, AURORA demonstrates even greater advantages on the challenging unseen test split. This consistent improvement on unseen categories indicates that AURORA possesses enhanced generalization capabilities, enabling it to effectively locate and segment novel objects not present in the training data. Figure 3 provides visual comparisons among the top-performing models.

### v2

@@ 1 Introduction
Humans perceive and interact with a complex, dynamic world by seamlessly integrating information from multiple sensory modalities. While vision plays a dominant role, auditory and textual cues are often crucial for accurately locating and comprehending objects of interest. This inherent multimodal nature of our perception highlights the necessity of developing systems that can achieve similar integrated understanding. Recent advancements in multimodal large language models, such as Qwen-Omni (xu2025qwen2), and VideoLLaMA2 (cheng2024videollama), have demonstrated significant progress in audio-visual comprehension. However, tasks like Reference Audio-Visual Segmentation (Ref-AVS) (wang2024ref) still remain challenging, requiring models to precisely segment specific sounding objects by integrating textual references, audio cues, and visual information. This demands sophisticated multimodal reasoning alongside accurate spatial localization capabilities.
@@
While previous works have explored fusion methods (wang2024ref; wang2025sam2; radman2025tsam) to align modalities, often employing an auxiliary text encoder with self-attention fusion, it remains unclear whether these models truly grasp the meaning of textual references or merely exhibit language bias. The “black box” nature of these approaches hinders genuine semantic understanding. To address these interpretability concerns, recent approaches have begun incorporating Large Language Models (LLMs) to enhance reasoning capabilities in Ref-AVS tasks, as demonstrated by Crab (du2025crab). However, their existing implementations often rely on supervised fine-tuning with pre-defined simple reasoning templates. This approach tends to lead the LLM towards “memorizing” these fixed patterns, which could result in outputs that reflect post-hoc rationalization rather than genuine and systematic reasoning. Moreover, the design of jointly training complex language reasoning with precise visual segmentation introduces new challenges, which risks compromising the pixel-level precision of the original segmentation models as pointed out by (liu2025seg).
@@
In this paper, we propose AURORA, a novel method that enhances both language comprehension and reasoning capabilities for audio-visual segmentation. Specifically, we leverage VideoLLaMA2 (cheng2024videollama) as our MLLM component to work in tandem with SAM (kirillov2023segment) for segmentation, enabling reasoning-guided segmentation. To prevent models from merely rationalizing the final output, we introduce a structured Chain-of-Thought (CoT) (wei2022chain) prompting mechanism via open-source Qwen-Omni (xu2025qwen2) to cost-effectively generate diverse reasoning paths for training. It involves distinct analytical steps focusing sequentially on visual, audio, and textual reference cues, followed by a final integration of these analyses to derive the answer. To mitigate the potential conflict between reasoning and segmentation during joint training, we introduce a segmentation feature distillation loss. This loss distills knowledge from a segmentation-only model into the joint model, enabling it to acquire CoT reasoning capabilities without compromising segmentation performance. To endow the model with genuine reasoning ability, we further adopt a two-stage training strategy after supervised fine-tuning (SFT) with CoT, combining reflective learning and reinforcement learning via Group Reward Policy Optimization (GRPO) (guo2025deepseek). In the corrective reflective-style stage, we transform low-quality reasoning paths from SFT-trained model into “reflective reasoning paths” by juxtaposing the initial flawed reasoning with a self-correction prompt followed by a Gemini-assisted high-quality revision. In the GRPO stage, we design a hybrid reward function incorporating format reward, IoU reward, and class reward to guide the model toward more robust and multimodally grounded reasoning.
@@ 2.1 Audio-Visual Segmentation
Audio-Visual Segmentation (AVS) aims to generate pixel-level masks for sounding objects and has attracted increasing research attention (zhou2023audio). While some prior works have incorporated text modality into AVS, they primarily focus on class alignment without fully exploiting the semantic richness of textual information (bhosale2023leveraging; bhosale2024unsupervised; luo2025tavis). Furthermore, traditional AVS methods (zhou2023audio; li2023catr; gao2024avsegformer; li2024qdformer; yang2024cooperation; ling2024transavs) lack explicit guidance mechanisms, making it challenging to identify specific objects of interest within complex audio-visual scenes. To address this limitation, Ref-AVS (wang2024ref) introduces a reference-based framework that provides continuous segmentation guidance through audio-visual-text fusion via a multi-modal cue aggregation module. Building upon this approach, SAM2-LOVE (wang2025sam2) further enhances the framework by integrating SAM2 with multimodal fusion, token propagation, and accumulation strategies. TSAM (radman2025tsam) enhances SAM with temporal modeling capabilities for spatio-temporal learning across video frames and replaces interactive prompting with data-driven prompts, thereby extending SAM’s functionality to dynamic video content. However, these methods treat text as an auxiliary modality and may fail to fully understand reference semantics.
Our work distinguishes itself by proposing a novel segmentation-centric framework integrating SAM (kirillov2023segment) with VideoLLaMA2 (cheng2024videollama) for reference-guided segmentation, employing CoT (wei2022chain) reasoning to incrementally decompose references and a segmentation feature distillation loss to preserve segmentation accuracy in SFT training.
@@ 2.2 Multimodal Language Models in Audio-Visual Scenes
In audio-visual understanding, predominant Multimodal Language Models (MLLMs) such as Qwen-Omni (xu2025qwen2), and VideoLLaMA2 (cheng2024videollama) integrate audio, visual, and textual modalities to achieve comprehensive scene understanding. MEERKAT (chowdhury2024meerkat) advances this field by constructing a fine-grained, large-scale audio-visual instruction-tuning dataset that endows models with temporal and spatial grounding capabilities. More specifically for reference AVS, CRAB (du2025crab) attempts to unify multimodal understanding and segmentation within an LLM framework. However, their reliance on supervised fine-tuning with pre-defined reasoning templates may lead to superficial pattern memorization rather than fostering genuine systematic reasoning or deep semantic comprehension, limiting their robustness in complex scenarios. To address these limitations, following SFT, we introduce an additional two-stage training process. The first stage, corrective reflective-style training, encourages the model to refine its prior knowledge and calibrate its foundational perceptual abilities. The second stage, GRPO, facilitates robust self-improvement by optimizing the model’s reasoning pathways based on targeted reward functions.
@@ 3 Methodology
In this work, we introduce AURORA, a novel reasoning-enhanced framework for reference audio-visual segmentation that addresses the fundamental challenge of achieving genuine multimodal reasoning and maintaining precise segmentation performance. We implement this framework by integrating SAM (kirillov2023segment) with VideoLLaMA2 (cheng2024videollama). Our approach employs a three-stage training procedure. In the first stage, we incorporate CoT reasoning using Qwen-Omni (xu2025qwen2) for SFT with segmentation feature distillation loss. In the second stage, we enhance the model’s foundational accuracy through a novel error-correction stage using structured “reflective examples” that address common perceptual and knowledge errors. The third stage employs GRPO (guo2025deepseek) to jointly enhance both reasoning capabilities and segmentation performance. Figure 1 illustrates the overall architecture of our proposed framework.
@@ 3.1 Stage1: SFT Training with CoT
Specifically, we first generate CoT answers using the open-source Qwen-Omni (xu2025qwen2) model to conserve computational costs compared to proprietary models, and establish the initial format through carefully designed prompts. Through prompt engineering, we decompose the reference reasoning into four distinct steps as shown below.
@@
where $D$ denotes the fine-tuning dataset, and $\pi_{\theta}$ is the MLLM parameterized by $\theta$. Since Ref-AVS is a binary segmentation task, we follow the LISA framework (lai2024lisa) and append the token “It is [SEG]” after the CoT reasoning. The hidden state feature of the [SEG] token is then used as a visual prompt for the SAM decoder.
@@
While CoT reasoning has demonstrated effectiveness in language tasks, it presents a challenge for multi-modal segmentation tasks. We hypothesize that fine-tuning with a joint objective may compromise the pixel precision of the segmentation models (liu2025seg), particularly when the language-based reasoning loss dominates the optimization dynamics. This, in turn, can degrade the representation of the [SEG] token, which is critical for triggering high-quality mask generation, thereby impairing segmentation performance. To address this issue and effectively decouple the optimization of reasoning and segmentation capabilities, we introduce a segmentation feature distillation loss as shown in the gray block in Figure 2.
@@ 3.2 Stage2: Corrective Reflective-Style Training
To endow our model with a more robust reasoning capability, we require high-quality examples of ideal thinking processes. We leverage a more powerful MLLM, Gemini, to serve as an expert annotator, whose role is to generate gold-standard reasoning paths by correcting flawed CoT samples. Specifically, we first prompt Gemini to critique the reasoning generated by our Stage 1 model. This process is focused on challenging samples from the training set, which we identify as those with both an IoU score below 0.6 and incorrect reasoning. This strict criterion ensures that corrective training targets only clear-cut failures where both the reasoning process and segmentation outcome are flawed. Adhering to the minimal modification principle (yu2024rlhf), Gemini is instructed to rectify any identified flaws—such as incorrect conclusions or superficial analysis of audio information—by making the edits (e.g., modifying, adding, or deleting words). We denote the initial, potentially flawed CoT from the Stage 1 model as $y_{wrong}$ and the corrected version from Gemini as $y_{correct}$. We then construct a “reflective path” as follows:
@@
$\mathcal{J\uparrow}$ $\mathcal{F\uparrow}$ $\mathcal{J}\!\&\!\mathcal{F}\uparrow$ $\mathcal{J\uparrow}$ $\mathcal{F\uparrow}$ $\mathcal{J}\!\&\!\mathcal{F\uparrow}$ $\mathcal{J\uparrow}$ $\mathcal{F}\uparrow$ $\mathcal{J}\!\&\!\mathcal{F}\uparrow$
@@
AVSBench (zhou2022audio) 23.2 51.1 37.2 32.4 54.7 43.5 27.8 52.9 40.3
AVSegFormer (gao2024avsegformer) 33.5 47.0 40.2 36.1 50.1 43.1 34.8 48.6 41.7
GAVS (wang2024prompting) 28.9 49.8 39.4 29.8 49.7 39.8 29.4 49.8 39.6
@@
EEMC (wang2024ref) 34.2 51.3 42.8 49.5 64.8 57.2 41.9 58.1 50.0
SAM-LAVS (wang2025sam2) 43.5 51.9 47.7 66.5 72.3 69.4 55.0 62.1 58.5
TSAM (radman2025tsam) 43.4 56.8 50.1 54.6 66.4 60.5 49.0 61.6 55.3
@@ 3.3 Stage3: Reinforcement Learning Training
Following the SFT stage, we incorporate reinforcement learning (RL) to further enhance the reasoning capability of the MLLM. Specifically, we adopt the GRPO (guo2025deepseek) framework to enable self-refinement based on the relative quality of generated outputs within a group. However, GRPO was originally designed for text generation and does not directly align with the segmentation nature of Ref-AVS. To address this, we redesign the reward function tailored for segmentation as shown in the green block of Figure 2.
@@
where ${r}_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\text{old}}(o_{i,t}|q,o_{i,<t})}$. Following (yu2025dapo), we set the KL divergence coefficient $\beta$ to zero, effectively removing this term for simplicity.
@@ 4.1 Experimental Settings
We evaluated our method on the Ref-AVS benchmark dataset (wang2024ref), which contains 4,000 videos with manual pixel-level annotations and expressions. The dataset is divided into a training set (2,908 videos), a validation set (276 videos), and a test set (818 videos). The test set is further split into three subsets: (1) seen split: containing the same 39 categories as the training set; (2) unseen split: with 13 additional categories not present during training; (3) null split: testing cases where expressions refer to non-existent or invisible objects. As in (wang2025sam2), we omit the null subset from evaluation.
@@
Our training unfolds in three stages. First, we perform an initial SFT for 720 steps using CoT reasoning generated by Qwen-Omni (xu2025qwen2). In the second stage, we conduct a 300-step corrective fine-tuning phase. For this, we curate 1,505 reflective-style examples by using Gemini 2.5-Pro to refine incorrect outputs from the first stage. These are then mixed with 3,500 standard CoT samples to prevent catastrophic forgetting. The final stage consists of 500 steps of GRPO (guo2025deepseek) training, where we generate 3 candidate responses per input for preference alignment. We use a batch size of 2 with 1 sample per device and gradient accumulation steps of 4. All experiments are conducted on two A40 GPUs.
@@
In line with the evaluation protocol of Ref-AVS (wang2024ref), we employ the Jaccard index ($\mathcal{J}$), F-score ($\mathcal{F}$), and their average ($\mathcal{J}\&\mathcal{F}$) as the primary evaluation metrics.
@@ 4.2 Main Results
We conduct a comprehensive performance comparison of our proposed AURORA against a suite of top-performing SOTA methods on the Ref-AVS benchmark. These include three audio-based segmentation methods (zhou2022audio; gao2024avsegformer; wang2024prompting), two visual-based segmentation methods (wu2022language; li2023robust), three multi-modal methods (wang2024ref; wang2025sam2; radman2025tsam), and one foundation-based method (du2025crab). The detailed results, presented in Table 1, demonstrate that AURORA achieves new state-of-the-art performance across multiple key metrics.
On the seen test split, our model surpasses the second-best performing method by substantial margins. We attribute this significant leap in performance to our framework’s core design: unlike prior methods that rely on simple modality fusion (wang2024ref; wang2025sam2; radman2025tsam) or train segmentation models from scratch (du2025crab), AURORA leverages the power of a large language model for deep semantic reasoning and integrates it with a pre-trained foundation model for segmentation. More importantly, AURORA demonstrates even greater advantages on the challenging unseen test split. This consistent improvement on unseen categories indicates that AURORA possesses enhanced generalization capabilities, enabling it to effectively locate and segment novel objects not present in the training data. Figure 3 provides visual comparisons among the top-performing models.